OpenStudy (anonymous):

Last week, Ethan finished 3 1⁄5 books, but this week he only finished 1 4⁄5 books. How many more books did he read last week? books

OpenStudy (anonymous):

So we want to subtract the second number by the first because, 1st is the no. of books last week, 2nd is this weeks, we want to work out the difference of the two numbers, this can also be known as subtract \[3\left(\begin{matrix}1 \\ 5\end{matrix}\right)\] - \[1 \left(\begin{matrix}4 \\ 5\end{matrix}\right)\]

OpenStudy (anonymous):

do the full numbers first resulting in 2 1/5

OpenStudy (anonymous):

so what do i do than can u helpme with my math. i might have more that i need help with

OpenStudy (anonymous):

you want to now see how to get from 2 1/5 to 4/5 which is minus 2/5

OpenStudy (anonymous):

thus you are left with 1/ 2/5

OpenStudy (anonymous):

another quicker way to do it is

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

put the whole numbers in the fraction as well last week books= 16/5 this weeks =9/5 now you just need to minus the numerators 16-5=7, therefore the answer is 7/5 you can turn that into the whole number again you will get 1 2/5

OpenStudy (anonymous):

answer is 7/5 or 1 2/5

OpenStudy (anonymous):

what else do you need help with?